OpenStudy (anonymous):

After the video 3, there's a homework which requires a few formulas from lecture 4. Is it supposed to be this way?

OpenStudy (anonymous):

I mailed OCW and this is their reply: The materials for 3.091SC have been distributed into separate modules and the content for each module has been developed to match the related topics as closely as possible. Yes, it is supposed to be structured this way. You can follow the materials in an order which is most convenient for you.

OpenStudy (anonymous):

The lecture notes seem to kind of cover the formula for the Bohr radius, but leave out Z (it is 1 for hydrogen).\[r_{n}=n^{2}h^{2} \epsilon _{o}/Z \pi m_{e}e^{2}; Z=1\] Wikipedia gives the formula for the Bohr radius...\[r_{n}=n^{2}ℏ^{2}/Z k_{e} e^{2} m_{e}\]...but, the Coulomb constant (not on list of fundamental physical constants)...\[k_{e}=1/4 \pi \epsilon _{o}\]...and ℏ (h-bar; listed fundamental constant)...\[ℏ=h/2 \pi\]...need to be substituted into the Wikipedia formula to get to the formula listed in the OCW materials. The formulae are out there, you just need to dig for them (for better and for worse).